英語 (English)

発信型英語 (Communicative English) (医 ((放 A)2 年))

座喜純 · 准教授/ 大学院ソシオ・アーツ・アンド・サイエンス研究部 2 単位前期月 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 This Communicative English class planned to give a good assistance

in rising up English speaking and communicating abilities. 【授業の概要】Classes will be running in ENGLISH language only for more

practices chance. There is no specific text book but there will be some other media depending on the ’Topics’ which will be selected up to date. Speaking out practices will take place during the classes.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】To refine listening skills and develop confidence in interpersonal

communication. 【授業の計画】

1. General Guidance

2. Selected Conversation 1

3. Selected Conversation 2

4. Selected Conversation 3

5. Selected Conversation 4

6. Selected Conversation 5

7. South Australia

8. Selected Conversation 6

9. Selected Conversation 7

10. Selected Conversation 8

11. Selected Conversation 9

12. Selected Conversation 10

13. Selected Conversation 11

14. Selected Conversation 12

15. Selected Conversation 13

16. Review 【教科書】 No specific text book but there will be some other media depending on

the ’Topic’ which will be selected up to date.

【参考書等】 [参考資料]

【成績評価の方法】 Attendance, Homework fulfillment and Participation in class.

【再試験の有無】 There is no make-up examination for this course.

【受講者へのメッセージ】 Punctuality and regular attendance are extremely important. Students are expected to communicate actively with all members of the class in English

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220893 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】

⇒ 座喜 . (オフィスアワー: contact E-mail: jzaki@ias.tokushima-u.ac.jp)

English

Communicative English (医 ((放 A)2 年))

Jun Zaki · Associate Professor / Institute of Socio-Arts and Sciences 2 units 前期月 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ This Communicative English class planned to give a good assistance in

rising up English speaking and communicating abilities. Outline⟩ Classes will be running in ENGLISH language only for more practices

chance. There is no specific text book but there will be some other media depending on the ’Topics’ which will be selected up to date. Speaking out practices will take place during the classes.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ To refine listening skills and develop confidence in interpersonal communi-

cation.

Schedule⟩

1. General Guidance

2. Selected Conversation 1

3. Selected Conversation 2

4. Selected Conversation 3

5. Selected Conversation 4

6. Selected Conversation 5

7. South Australia

8. Selected Conversation 6

9. Selected Conversation 7

10. Selected Conversation 8

11. Selected Conversation 9

12. Selected Conversation 10

13. Selected Conversation 11

14. Selected Conversation 12

15. Selected Conversation 13

16. Review Textbook⟩ No specific text book but there will be some other media depending

on the ’Topic’ which will be selected up to date.

Reference⟩ [参考資料]

Evaluation Criteria⟩ Attendance, Homework fulfillment and Participation in class. Re-evaluation⟩ There is no make-up examination for this course. Message⟩ Punctuality and regular attendance are extremely important. Students are expected to communicate actively with all members of the class in English

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220893 Contact (Office-Hour, Room, E-mail)⟩

⇒ Zaki . (Office Hour: contact E-mail: jzaki@ias.tokushima-u.ac.jp)